Use the following table to answer questions 10 to 12. Type of School Location Public (P Religious (R) Other Private (P Row Total 15 10 70 80 35 Inner City (D 35 25 Suburban (S 45 25 Rural (R 50 185 Col Total 105 30 10. What is the probability that a student attends a public school and is in a rural area? 11. Given a student attends a religious school, what is the probability he/she is in a rural location? 12. If a student lives in an inner city location, what is the probability that he/she is in public school?
Answer #1

10.

P(Student attending a public school | Student is from a rural area)

= No. of students attending a public school in a rural area/No. of students in the rural area

= 25\35

=5\7

11.

Let's take Religious = Re

P(R|Re) = P(R\cap Re)\div P(Re)

= (5/185)/(30/185)

= 1/6

12.  

P(P|I) = P(P\cap I)\div P(I)

= (35/185)/(70/185)

=1/2